OpenStudy (anonymous):

A container that holds the water for the football team is 1/2 full. After pouring 8 gallons of water, it is 5/6 full. How many gallons can the container hold?

OpenStudy (anonymous):

if x is the capacity in gallons x/2 + 8 = 5x/6 => x=24

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

the difference between being 1/2 full and 5/6 full is 8 gallons. subtract 1/2 from 5/6 to find that 8 gallons=1/3 capacity. multiply 8 gallons by the reciprocal of 1/3 (which is 3) to find the total capacity of the container. 24 Gallons